OpenStudy (anonymous):

how would you factor 8r^6+27s^12

OpenStudy (cwrw238):

make use of the identity a^3 + b^3 = (a + b)(a^2 - ab + b^2) hold on- i'll check to see if i remember that one correctly)

OpenStudy (cwrw238):

yes thats right so if you compare your expression with the above you get a = 2r^2 ( because 8r^6 = (2r^2)^3 and b = 3s^4 so plug these in to the formula for a^3 + b^3 and you've got it